A truck on a straight road starts from rest, accelerating at 2.00 m/s2 until it reaches a speed of 20.0 m/s. Then the truck travels for 20 s at a constant speed until the brakes are applied, stopping the truck in a uniform manner in an additional 5.00 s. (a) How long is the truck in motion? (b) What is the average velocity of the truck for the motion described?
1
Expert's answer
2020-10-07T10:13:04-0400

a) Since the speed is "V = a\times t" , then "t_1=\frac{V}{a_1}" , where "a_1" is the acceleration in the first interval, "t_1" is the time spent on the first interval, then

"t_1=\frac {20} 2=10" s.

By condition "t_2=20" c, and "t_3=5" c, therefore, the total time is "t = 10 + 20 + 5 = 35" s.

b) Average speed is determined by formula "V_{ave}=\frac {S_{total}}{t_{total}}" ;

The path for the first interval is "S_1=\frac {a_1\times t_1^2}{2}=\frac {2\times 10 ^ 2} {2} = 100" m;

For the second interval, the path is equal to "S_2=V\times t_2= 20\times 20 = 400" m;

And for the third "S_3=\frac {V_1^2-V_0^2}{2\times a_2}" , where "V_1=0" , since this is the final speed at the stop, "V_0" is the initial speed at the third interval, "a_2" is the acceleration at the third interval and it is equal to "a_2=\frac {V_0}{t_3}=\frac {20} {5} = 4" m/s "^ 2" , since the acceleration is directed against the velocity vector then "a_2=-4" m/s "^ 2"

"S_3=\frac {0-20^2}{-4\times 2}=50" m;

As a result, we have:

"V_{ave}=\frac {400+100+50}{35}\approx15.71" m/s.